QUANDALEDINGL3:

Solve for x and show your steps. Is the solution extraneous? Check your work to show how you determined if the solution is extraneous or not. sqrt 2x minus 7 equal 1

dude:

Hmm do you mean \(\sqrt{2x}-7=1\) or \(\sqrt{2x-7}=1\)?

QUANDALEDINGL3:

thanks

QUANDALEDINGL3:

yep